### v1

@@ 4.4 Simulation specifications for models on complete graphs
We first simulate our topic-weighted BCMs on complete graphs. We simulate our topic-weighted HK model on a 2000-node complete graph. Our topic-weighted DW model is computationally more expensive than our topic-weighted HK model; we simulate it on a 500-node complete graph. In our simulations, we use the initial opinion distributions described in section 4.3. That is, we examine uniform, wedge, $G(0.22,0)$, and $G(0.22,0.8)$ initial opinions.
We consider values of the confidence bound $c$ that allow us to examine the transition between opinion consensus and fragmentation. For our topic-weighted HK model, we use $c\in[0.025,0.3]$ in our simulations. For our topic-weighted DW model, we use $c\in[0.05,0.4]$ in our simulations. For both topic-weighted BCMs, we use $\omega\in[0.1,1]$ in our simulations.
In our simulations, both the initial opinions and the selection of interacting node pairs and topics are random. To mitigate the impact of randomness, we repeat the simulations and compute averaged quantities. For our topic-weighted HK model, we randomly generate 25 sets of initial opinions for each initial opinion distribution. We reuse those sets of initial opinions in our simulations with different values of the BCM parameters (namely, the confidence bound $c$ and the topic weight $\omega$). Similarly, for our topic-weighted DW model, we randomly generate 25 sets of initial opinions for each initial opinion distribution.
@@ 4.5 Simulation specifications for our topic-weighted HK model on stochastic-block-model (SBM) graphs
We use numerical simulations to investigate how the network community structure and community-dependent initial opinions affect opinion evolution. To do this, we simulate our topic-weighted HK model on undirected two-community stochastic-block-model (SBM) networks [31, 43]. We do not run analogous simulations for our topic-weighted DW model because the model has significantly longer simulation times than our topic-weighted HK model.
@@
For our simulations with community-dependent initial opinions, nodes in community A have initial opinions that lie in $[0,0.5]\times[0,1]$, and nodes in the community B have initial opinions that lie in $[0.5,1]\times[0,1]$. Thus, the two communities initially have split views on topic 1, but not on topic 2. For the community-dependent uniform and wedge distributions, we select the initial opinions for topic 1 uniformly at random from $[0,0.5]$ for nodes in community A and from $[0.5,1]$ for nodes in community B. We then determine the corresponding opinions for topic 2 by uniformly random selection from $[0,1]$ for the uniform distribution and by computing $x^{2}=1-|2x^{1}-1|$ for the wedge distribution. For the truncated Gaussian $G(0.22,0.8)$ distribution, we start by drawing 2000 opinions, one for each node in our SBM graph. We reject individual opinions and resample until we have exactly 1000 opinions that lie in $[0,0.5]\times[0,1]$ for nodes in community A and 1000 opinions that lie in $[0.5,1]\times[0,1]$ for nodes in community B.
For each of our simulations of our topic-weighted HK model with community-dependent initial opinions, we also run a corresponding simulation with community-independent initial opinions. This corresponding simulation uses the same two-community SBM graph, set of initial opinions, and BCM parameters. However, we shuffle the set of initial opinions and randomly assign them to nodes in the graph. These corresponding simulations give us a fair comparison that allows us to investigate the effect of having community-dependent initial opinions without introducing effects from varying the graph or set of initial opinions.
We randomly generate 5 two-community SBM graphs and 10 sets of initial opinions for each random graph. We reuse these 50 combinations of the random graph and set of initial opinions in our simulations with different values of the BCM parameters (namely, the confidence bound $c$ and the topic weight $\omega$). We use BCM parameter values $c\in[0.025,0.5]$ and $\omega\in[0.1,1]$ in our simulations.
@@ 5.1.1 A complete graph
In our simulations, the choice of initial opinion distribution significantly influences whether our topic-weighted HK model converges to a consensus state or a fragmented state. For fixed $\omega$, the transition between consensus and fragmentation occurs at smaller values of $c$ for the two Gaussian initial opinion distributions than for the uniform and wedge initial opinion distributions. Intuitively, it makes sense that the two Gaussian distributions promote less fragmentation than the uniform and wedge distributions. This is because the initial opinions for the Gaussian distributions are more concentrated near the center (see Fig. 4) and less concentrated near the boundary of the opinion space $[0,1]\times[0,1]$. For the uniform and Gaussian $G(0.22,0)$ distributions, as we decrease $\omega$, the value of $c$ at which the transition between consensus and fragmentation occurs also decreases. For the uniform distribution and $c=0.2$, our topic-weighted HK model with $\omega\leq 0.4$ reaches consensus, while the baseline HK model has opinion fragmentation.
@@
To investigate the impact of the distribution of initial opinions on the distribution of final opinions, we plot the trajectories of the opinions of nodes in our simulations. In our opinion space $[0,1]\times[0,1]$, for each node, we plot the path of its opinion; that is, we plot how its opinion changes from its initial opinion to its opinion at convergence time $T_{f}$. In Fig. 6, we show some plots of the opinion trajectories for simulations with uniform initial opinions and $c=0.1$ for both our topic-weighted HK model with $\omega=0.1$ and the baseline HK model. In the simulation of the baseline HK model (see Fig. 6B), for each final opinion cluster, the initial opinions of the nodes in that opinion cluster appear to visually form a rectangle containing the corresponding the mean final opinion; the opinion space can be split into non-overlapping rectangular regions that cover the space and nodes with initial opinions in the same rectangular region are in the same final opinion cluster. By contrast, in the simulation of our topic-weighted HK model with $\omega=0.1$ (see Fig. 6A), for each final opinion cluster, the initial opinions of nodes that opinion cluster do not form rectangular shapes. The simulation of our topic-weighted HK model with $\omega=0.1$ has $Q(T_{f})\approx 0.1956$ and there is less opinion fragmentation than the simulation of the baseline HK model, which has $Q(T_{f})\approx 0.0665$. Recall (see Fig. 2) that in 2D, the baseline HK model has square regions of receptiveness and our topic-weighted HK model with $\omega\in(0,1)$ has 8-sided regions of receptiveness. For fixed $c$, as we decrease $\omega$, the area of the regions of receptiveness increases. We hypothesize that the square shape of the regions of receptiveness for the baseline HK model yields the rectangular regions in Fig. 6B. The larger size and non-square shape of the regions of receptiveness for our topic-weighted HK model with $\omega=0.1$ result in nodes with initial opinions in larger and more irregular regions of opinion space being pulled into the same final opinion cluster, yielding less opinion fragmentation than the baseline model. We hypothesize that a similar effect applies to the truncated Gaussian distributions, causing smaller values of $\omega$ in our topic-weighted HK model to promote less opinion fragmentation. We show some plots of opinion trajectories for the truncated Gaussian distributions of initial opinions $G(0.22,0)$ (see Fig. 13) and $G(0.22,0.8)$ (see Fig. 14) in section A.1.
@@ 5.1.2 Two-community SBM graphs
We now discuss the simulations of our topic-weighted HK model on two-community SBM random graphs. As we discussed in section 4.5, our SBM random graphs consist of two communities with 1000 nodes each. For our simulations, we generate five random SBM graphs. For each graph and each initial opinion distribution, we use 10 distinct sets of initial opinions in Monte Carlo simulations of our topic-weighted HK model. For each combination of random SBM graph, set of initial opinions, and set of BCM parameters (namely, the confidence bound $c$ and topic weight $\omega$), we simulate two scenarios: one with community-dependent assignment of initial opinions and one with random assignment of initial opinions.
For each of the uniform, wedge, and truncated Gaussian $G(0.22,0)$ initial opinion distributions and fixed BCM parameters (namely, $c$ and $\omega$), when initial opinions are randomly assigned to nodes, the final order parameter $Q(T_{f})$ (see (70)) for the two-community SBM graphs is similar to $Q(T_{f})$ for a complete graph. The minimum value of $c$ for our simulations to always reach consensus is larger when we assign initial opinions to nodes based on their community compared to when we randomly assign initial opinions to nodes. Additionally, for the community-dependent initial opinions, as we decrease $\omega$, the value of $c$ required for simulations to reach consensus also decreases. We observe that when consensus is not reached and the initial opinions are assigned based on node community, the nodes in a final opinion cluster all belong to the same community. It appears that the two-community structure of our SBM graphs has little effect on the amount of opinion fragmentation when initial opinions are independent of community, but tends to increase opinion fragmentation when we have community-dependent initial opinions in which the communities are “split” on their initial opinions on one topic (specifically, nodes in one community have $x^{1}(0)\in[0,0.5]$ and nodes in the other community have $x^{1}(0)\in[0.5,1]$; see section 4.5).
In Fig. 8, we show the final order parameters $Q(T_{f})$ (see (70)) of our simulations of our topic-weighted HK model on two-community SBM graphs with uniform initial opinions. For uniform initial opinions that are randomly assigned to nodes (see Fig. 8B) and fixed BCM parameters, $Q(T_{f})$ is similar to what we observe for a complete graph. For a uniform initial opinion distribution, our simulations on SBM graphs, both those with community-dependent initial opinions and those with randomly-assigned initial opinions, have the same general trends as our simulations on complete graph; namely, we tend to observe more opinion fragmentation as we either (1) decrease $c$ for fixed $\omega$ or (2) decrease $\omega$ for fixed $c$. For fixed $\omega$, a larger value of $c$ is required for simulations to always reach consensus for the community-dependent initial opinions (see Fig. 8A) than for randomly-assigned initial opinions (see Fig. 8B).
@@
For all our simulations with community-dependent initial opinions, when consensus is not reached, we observe that nodes in the same final opinion cluster also belong to the same community. In Fig. 9, we show the opinion trajectories for two simulations of our topic-weighted HK model with $\omega=0.1$ and $c=0.1$ and uniform initial opinions. In Fig. 9A, each community can be split into three corresponding final opinion clusters. By contrast (see Fig. 9B), when initial opinions are randomly assigned to nodes, final opinion clusters can consist of nodes from each community. For community-dependent initial opinions and intermediate values of $c$, we observe that $Q(T_{f})=0.5$ (see Fig. 8). In this situation, there are two final opinion clusters; each final opinion cluster contains all the nodes from a single community.
@@
In Fig. 10, we show the final order parameters $Q(T_{f})$ of our simulations of our topic-weighted HK model on two-community SBM graphs with wedge initial opinions. Unlike for the community-dependent uniform and $G(0.22,0)$ initial opinions, for the community-dependent wedge initial opinions, for small $\omega$, as we increase $c$, the final order parameter $Q(T_{f})$ does not always increase or stay the same. For $\omega\in\{0.1,0.2,0.3\}$, we observe that $Q(T_{f})$ when $c=0.2$ is smaller than $Q(T_{f})$ when either $c=0.175$ or $c=0.225$. In Fig. 11, we show two opinion trajectory plots for simulations of our topic-weighted HK model with $\omega=0.1$. Both simulations use the same community-dependent wedge initial opinions. In both simulations, the mean opinions associated with the two final opinion clusters are just outside of each other’s regions of receptiveness. One can imagine that small perturbations of the initial opinions could result in a simulation that reaches consensus. In the two simulations shown, the opinions of the two final opinion clusters are closer to each other’s region of absorption than the opinions of the two final opinion clusters for the simulation with $c=0.2$ (see Fig. 11A) than for the simulation with $c=0.225$ (see Fig. 11B). Therefore, it makes sense that when $\omega\leq 0.3$, we observe that more simulations of our topic-weighted HK model with $c=0.2$ reach consensus than simulations of our model with $c=0.225$. We hypothesize that we only observe this behavior for the wedge initial opinions and not the uniform or Gaussian initial opinions because the shape of the wedge distribution causes the mean opinions of the final opinion clusters to be close to each other’s regions of absorption when there are two final opinion clusters.
@@ 5.2 Topic-weighted DW model
In Fig. 12, we show the final order parameters $Q(T_{f})$ (see (70)) in our simulations of our topic-weighted DW model for various initial opinion distributions. Overall, we observe similar trends for $Q(T_{f})$ for our topic-weighted DW model as we did for our topic-weighted HK model. In this section, we discuss our observations of $Q(T_{f})$ for our topic-weighted DW model. In Appendix B, we show and discuss some plots of opinion trajectories in our topic-weighted DW model.
For the uniform and Gaussian initial opinion distributions, we observe similar trends for our topic-weighted DW model as we did for our topic-weighted HK model. One trend is that, for fixed confidence bound $c$ and a fixed initial opinion distribution, we tend to observe less opinion fragmentation as we decrease the topic weight $\omega$. Another trend is that, as we decrease $\omega$, we observe a decrease in the value of $c$ at which the transition between consensus and fragmentation occurs. For our topic-weighted DW model with fixed $\omega$ and fixed $c$, the Gaussian initial opinion distributions tend to have less opinion fragmentation that the uniform initial opinion distribution. However, for fixed $\omega$, the value of $c$ above which all simulations reach a consensus (i.e., $Q(T_{f})=1$ for all simulations) appears similar for the uniform and Gaussian initial opinion distributions. For our simulations of our topic-weighted DW model, the transition between consensus and fragmentation is not as clear as for our simulations of our topic-weighted HK model (see Fig. 5). We hypothesize that finite-size effects from the 500-node complete graph and randomness from selecting pairs of nodes to interact for our topic-weighted DW simulations may contribute to this observation.

### v2

@@ 4.4 Simulation specifications for models on complete graphs
We first simulate our topic-weighted BCMs on complete graphs. We simulate our topic-weighted HK model on a 2000-node complete graph. Our topic-weighted DW model is computationally more expensive than our topic-weighted HK model; we simulate it on a 500-node complete graph. In our simulations, we examine uniform, wedge, $G(0.22,0)$, and $G(0.22,0.8)$ initial opinions.
We consider values of the confidence bound $c$ that allow us to examine the transition between opinion consensus and fragmentation. We consider $c\in[0.025,0.3]$ and $c\in[0.05,0.4]$ for our topic-weighted HK simulations and our topic-weighted DW simulations, respectively. For both topic-weighted BCMs, we use $\omega\in[0.1,1]$ in our simulations.
In our simulations, both the initial opinions and the selection of interacting node pairs and topics are random. To mitigate the impact of randomness, we repeat the simulations and compute averaged quantities. We randomly generate 25 sets of initial opinions for each initial opinion distribution. We reuse those sets of initial opinions in our simulations with different values of the BCM parameters (namely, $c$ and $\omega$).
@@ 4.5 Simulation specifications for our topic-weighted HK model on stochastic-block-model (SBM) graphs
We use numerical simulations to investigate how the network community structure and community-dependent initial opinions affect opinion evolution. To do this, we simulate our topic-weighted HK model on undirected two-community stochastic-block-model (SBM) networks [32, 44]. We do not run analogous simulations for our topic-weighted DW model because of its long simulation times.
@@
For our simulations with community-dependent initial opinions, nodes in community A have initial opinions that lie in $0,0.5]\times[0,1]$, and nodes in the community B have initial opinions that lie in $0.5,1]\times[0,1]$. Thus, the two communities initially have split views on topic 1, but not on topic 2. For the community-dependent uniform and wedge distributions, we select the initial opinions for topic 1 uniformly at random from $[0,0.5]$ for nodes in community A and from $[0.5,1]$ for nodes in community B. We then determine the corresponding opinions for topic 2 by uniformly random selection from $[0,1]$ for the uniform distribution and by computing $x^{2}=1-|2x^{1}-1|$ for the wedge distribution. For the truncated Gaussian $G(0.22,0.8)$ distribution, we start by drawing 2000 opinions, one for each node in our SBM graph. We reject individual opinions and resample until we have exactly 1000 opinions that lie in $0,0.5]\times[0,1]$ for nodes in community A and 1000 opinions that lie in $0.5,1]\times[0,1]$ for nodes in community B.
For each of our simulations of our topic-weighted HK model with community-dependent initial opinions, we also run a corresponding simulation with community-independent initial opinions. This corresponding simulation uses the same two-community SBM graph, set of initial opinions, and BCM parameters. However, we shuffle the set of initial opinions and randomly assign them to nodes in the graph. These simulations allow us to investigate the effect of having community-dependent initial opinions without introducing effects from varying the graph or set of initial opinions.
We randomly generate 5 two-community SBM graphs and 10 sets of initial opinions for each random graph. We reuse these 50 combinations of the random graph and set of initial opinions in our simulations with different values of the BCM parameters (namely, $c$ and $\omega$). We use BCM parameter values $c\in[0.025,0.5]$ and $\omega\in[0.1,1]$ in our simulations.
@@ 5.1.1 A complete graph
In our simulations, the choice of initial opinion distribution significantly influences whether our topic-weighted HK model converges to a consensus state or a fragmented state. For fixed $\omega$, the transition between consensus and fragmentation occurs at smaller values of $c$ for the two Gaussian initial opinion distributions than for the uniform and wedge initial opinion distributions. Intuitively, it makes sense that the two Gaussian distributions promote less fragmentation than the uniform and wedge distributions. This is because the initial opinions for the Gaussian distributions are more concentrated near the center (see Fig. 4) and less concentrated near the boundary of the opinion space $0,1]\times[0,1]$. For the uniform and Gaussian $G(0.22,0)$ distributions, as we decrease $\omega$, the value of $c$ at which the transition between consensus and fragmentation occurs also decreases. For the uniform distribution and $c=0.2$, our topic-weighted HK model with $\omega\leq 0.4$ reaches consensus, while the baseline HK model has opinion fragmentation.
@@
To investigate the impact of the distribution of initial opinions on the distribution of final opinions, we plot the trajectories of the opinions of nodes in our simulations. In our opinion space $0,1]\times[0,1]$, for each node, we plot the path of its opinion; that is, we plot how its opinion changes from its initial opinion to its opinion at convergence time $T_{f}$. In Fig. 6, we show some plots of the opinion trajectories for simulations with uniform initial opinions and $c=0.1$ for both our topic-weighted HK model with $\omega=0.1$ and the baseline HK model. In the simulation of the baseline HK model (see Fig. 6B), for each final opinion cluster, the initial opinions of the nodes in that opinion cluster appear to visually form a rectangle containing the corresponding the mean final opinion; the opinion space can be split into non-overlapping rectangular regions that cover the space and nodes with initial opinions in the same rectangular region are in the same final opinion cluster. By contrast, in the simulation of our topic-weighted HK model with $\omega=0.1$ (see Fig. 6A), for each final opinion cluster, the initial opinions of nodes in that opinion cluster do not form rectangular shapes. The simulation of our topic-weighted HK model with $\omega=0.1$ has $Q(T_{f})\approx 0.1956$ and there is less opinion fragmentation than the simulation of the baseline HK model, which has $Q(T_{f})\approx 0.0665$. Recall (see Fig. 2) that in 2D, the baseline HK model has square regions of receptiveness and our topic-weighted HK model with $\omega\in(0,1)$ has 8-sided regions of receptiveness. For fixed $c$, as we decrease $\omega$, the area of the regions of receptiveness increases. We hypothesize that the square shape of the regions of receptiveness for the baseline HK model yields the rectangular regions in Fig. 6B. The larger size and non-square shape of the regions of receptiveness for our topic-weighted HK model with $\omega=0.1$ result in nodes with initial opinions in larger and more irregular regions of opinion space being pulled into the same final opinion cluster, yielding less opinion fragmentation than the baseline model. We hypothesize that a similar effect applies to the truncated Gaussian distributions, causing smaller values of $\omega$ in our topic-weighted HK model to promote less opinion fragmentation. We show some plots of opinion trajectories for the truncated Gaussian distributions of initial opinions $G(0.22,0)$ (see Fig. 13) and $G(0.22,0.8)$ (see Fig. 14) in section A.1.
@@ 5.1.2 Two-community SBM graphs
We now discuss the simulations of our topic-weighted HK model on two-community SBM random graphs. As we discussed in section 4.5, our SBM random graphs consist of two communities with 1000 nodes each. For our simulations, we generate five random SBM graphs. For each graph and each initial opinion distribution, we use 10 distinct sets of initial opinions in Monte Carlo simulations of our topic-weighted HK model. For each combination of random SBM graph, set of initial opinions, and set of BCM parameters, we simulate two scenarios: one with community-dependent assignment of initial opinions and one with random assignment of initial opinions.
For each of the uniform, wedge, and truncated Gaussian $G(0.22,0)$ initial opinion distributions and fixed BCM parameters, when initial opinions are randomly assigned to nodes, the final order parameter $Q(T_{f})$ (see (70)) for the two-community SBM graphs is similar to $Q(T_{f})$ for a complete graph. The minimum value of $c$ for our simulations to always reach consensus is larger when we assign initial opinions to nodes based on their community, compared to when we randomly assign initial opinions to nodes. It appears that the two-community structure of our SBM graphs has little effect on the amount of opinion fragmentation when initial opinions are independent of community, but tends to increase opinion fragmentation when we have community-dependent initial opinions in which the communities are “split” on their initial opinions on one topic (specifically, nodes in one community have $x^{1}(0)\in[0,0.5]$ and nodes in the other community have $x^{1}(0)\in[0.5,1]$; see section 4.5). For all our simulations with community-dependent initial opinions, when consensus is not reached, nodes in the same final opinion cluster also belong to the same community.
In Fig. 8, we show the final order parameters $Q(T_{f})$ of our simulations of our topic-weighted HK model on two-community SBM graphs with uniform initial opinions. For uniform initial opinions, our simulations on SBM graphs, both those with community-dependent initial opinions and those with randomly-assigned initial opinions, have the same general trends as our simulations on complete graph; namely, we tend to observe more opinion fragmentation as we either (1) decrease $c$ for fixed $\omega$ or (2) increase $\omega$ for fixed $c$. For fixed $\omega$, a larger value of $c$ is required for simulations to always reach consensus for the community-dependent initial opinions (see Fig. 8A) than for randomly-assigned initial opinions (see Fig. 8B).
@@
In Fig. 9, we show the opinion trajectories for two simulations of our topic-weighted HK model with $\omega=0.1$ and $c=0.1$ and uniform initial opinions. In Fig. 9A, each final opinion cluster is contained within one community, and each community can be split into three corresponding final opinion clusters. By contrast (see Fig. 9B), when initial opinions are randomly assigned to nodes, final opinion clusters can consist of nodes from each community. For community-dependent initial opinions and intermediate values of $c$, we observe that $Q(T_{f})=0.5$ (see Fig. 8). In this situation, there are two final opinion clusters; each final opinion cluster contains all the nodes from a single community.
@@
In Fig. 10, we show the final order parameters $Q(T_{f})$ of our simulations of our topic-weighted HK model on two-community SBM graphs with wedge initial opinions. Unlike for the community-dependent uniform and $G(0.22,0)$ initial opinions, for the community-dependent wedge initial opinions, for small $\omega$, as we increase $c$, the final order parameter $Q(T_{f})$ does not always increase or stay the same. For $\omega\in\{0.1,0.2,0.3\}$, we observe that $Q(T_{f})$ when $c=0.2$ is smaller than $Q(T_{f})$ when either $c=0.175$ or $c=0.225$. In Fig. 11, we show two opinion trajectory plots for simulations of our topic-weighted HK model with $\omega=0.1$. Both simulations use the same community-dependent wedge initial opinions. In both simulations, the mean opinions associated with the two final opinion clusters are just outside of each other’s regions of receptiveness. One can imagine that small perturbations of the initial opinions could result in a simulation that reaches consensus. In the two simulations shown, the opinions of the two final opinion clusters are closer to each other’s region of absorption for the simulation with $c=0.2$ (see Fig. 11A) than for the simulation with $c=0.225$ (see Fig. 11B). Therefore, it makes sense that when $\omega\leq 0.3$, we observe that more simulations of our topic-weighted HK model with $c=0.2$ reach consensus than simulations of our model with $c=0.225$. We hypothesize that we only observe this behavior for the wedge initial opinions and not the uniform or Gaussian initial opinions because the shape of the wedge distribution causes the mean opinions of the final opinion clusters to be close to each other’s regions of absorption when there are two final opinion clusters.
@@ 5.2 Topic-weighted DW model
In Fig. 12, we show the final order parameters $Q(T_{f})$ (see (70)) in our simulations of our topic-weighted DW model for various initial opinion distributions. Overall, we observe similar trends for $Q(T_{f})$ for our topic-weighted DW model as we did for our topic-weighted HK model. In this section, we discuss our observations of $Q(T_{f})$ for our topic-weighted DW model. In Appendix B, we show and discuss some plots of opinion trajectories.
For the uniform and Gaussian initial opinion distributions, we observe similar trends for our topic-weighted DW model as we did for our topic-weighted HK model. One trend is that, for fixed confidence bound $c$ and a fixed initial opinion distribution, we tend to observe less opinion fragmentation as we decrease the topic weight $\omega$. Another trend is that, as we decrease $\omega$, we observe a decrease in the value of $c$ at which the transition between consensus and fragmentation occurs. For our topic-weighted DW model with fixed $\omega$ and fixed $c$, the Gaussian initial opinion distributions tend to have less opinion fragmentation than the uniform initial opinion distribution. However, for fixed $\omega$, the value of $c$ above which all simulations reach a consensus (i.e., $Q(T_{f})=1$ for all simulations) appears similar for the uniform and Gaussian initial opinion distributions. For our simulations of our topic-weighted DW model, the transition between consensus and fragmentation is not as clear as for our simulations of our topic-weighted HK model (see Fig. 5). We hypothesize that finite-size effects from the 500-node complete graph and randomness from selecting pairs of nodes to interact for our topic-weighted DW simulations may contribute to this observation.